# Generalization Below the Edge of Stability: The Role of Data Geometry

### v1

@@ Abstract
Understanding generalization in overparameterized neural networks hinges on the interplay between the data geometry, neural architecture, and training dynamics. In this paper, we theoretically explore how data geometry controls this implicit bias. This paper presents theoretical results for overparameterized two-layer ReLU networks trained below the edge of stability. First, for data distributions supported on a mixture of low-dimensional balls, we derive generalization bounds that provably adapt to the intrinsic dimension. Second, for a family of isotropic distributions that vary in how strongly probability mass concentrates toward the unit sphere, we derive a spectrum of bounds showing that rates deteriorate as the mass concentrates toward the sphere. These results instantiate a unifying principle: When the data is harder to “shatter” with respect to the activation thresholds of the ReLU neurons, gradient descent tends to learn representations that capture shared patterns and thus finds solutions that generalize well. On the other hand, for data that is easily shattered (e.g., data supported on the sphere) gradient descent favors memorization. Our theoretical results consolidate disparate empirical findings that have appeared in the literature.
@@ 1 Introduction
Contributions. In this work, we argue that the effectiveness of this data-dependent regularity is governed by a single, unifying principle, which we term data shatterability
The less shatterable the data geometry, the stronger the implicit regularization of EoS becomes.
We formalize this principle with the following theoretical results:
  • Provable Adaptation to Low-dimensionality. Assuming that the input features are supported only on a mixture of $m$-dimensional subspaces in $\mathbb{R}^{d}$ with $m<d$, we prove the generalization bound of $\tilde{O}(n^{-\frac{1}{2m+4}})$ that adapts to the intrinsic dimension as Theorem 3.2, where $\tilde{O}$ hides the constants that mildly depends on the number of subspaces (at most linear), and the logarithmic factor of the probability term. Our empirical results on synthetic data also support the theoretical analysis.
@@
  • A Spectrum of Generalization on Isotropic Data. We show that generalization performance smoothly degrades as data concentrates near the boundary of its support. We provide precise upper bounds that depend on the dimension $d$ and a concentration parameter $\alpha$ (Theorem 3.5), as well as lower bounds (Theorem 3.6). This analysis culminates in the limit of extreme boundary concentration (data on a sphere), where we provide a concrete construction of a network that perfectly interpolates any dataset at the BEoS regime, see Theorem 3.7. In particular, the “neural shattering” phenomenon, identified by Liang et al. (2025) for the uniform ball distribution, represents one special point of the broader generalization spectrum we uncover.
Our theoretical results for both subspace mixtures and isotropic distributions demonstrate how the principle of data shatterability operates in two distinct ways.
On the one hand, for data residing on a mixture of subspaces, its low-dimensional nature inherently limits its shatterability. We demonstrate that the network’s complex decision boundaries, formed by combinations of half-spaces, are fundamentally constrained by the data’s intrinsic low-dimensional structure. For example, when data lies on a line within $\mathbb{R}^{d}$, a ReLU’s complex hyperplane boundaries reduce to a series of knots and entire complexity is defined by the locations and magnitudes of these knots (Figure. 1a). The cornerstone of our proof is showing that the stability induced, data-dependent implicit regularization is adaptive to this nature.
On the other hand, in the isotropic case, shatterability is governed by the data’s radial concentration. As more mass concentrates towards the spherical shell, the data geometry becomes more shatterable. This is because the more the data is concentrated near the sphere, the more non-overlapping caps the network can create across the sphere, while keeping the mass of data inside each cap constant (see the comparison of Figure 1b & Figure 1c). This allows the network to partition the data into a large number of disjoint, sparsely populated regions, a key feature of high shatterability, and dedicate different neurons to memorizing the labels within each region. Our theoretical analysis and empirical observation confirms that such solutions can be dynamically stable, meaning they can be favored by gradient descent when the data geometry permits.
(a) ReLUs shatter lines by knots (b) More caps ($\alpha=0.5$) (c) Less caps ($\alpha=2$)
Figure 1: Visualization of Data Shatterability. (a) Knots as learned representations: When data lies on a mixture of lines (black lines), the decision boundaries of ReLU neurons (translucent planes) reduce to a finite set of knots (colored dots). (b & c) More caps v.s. Less caps: Red dots depict data points and dashed lines represent neuron boundaries that partition the data support into non-overlapping caps (shaded regions), each containing equal probability mass.
From a representation learning perspective, our constructions for the lower bound and flat interpolation (Theorems 3.6 and 3.7) demonstrate that data shatterability directly governs the model’s learning strategy. High shatterability enables the formation of “memorizing neurons” that activate on only a few examples [^1] For example, in Figure 1b, points in different caps can be mapped to orthogonal hidden representations, and a larger number of caps induces a higher effective dimension in the representation space, thus increasing the network’s expressive capacity to memorize the labels.. Conversely, low shatterability makes this memorization strategy difficult, implicitly forcing the optimizer to learn robust, shared representations (see Figure 3). We thus propose data shatterability as a foundational principle explaining the representations favored by gradient descent, bridging a critical gap between learning theory and practice.
@@
Related work and novelty. We build upon a recent line of work (Qiao et al., 2024; Liang et al., 2025) that theoretically study the generalization of neural networks in Edge-of-Stability regime (Cohen et al., 2020) from a function space perspective (Mulayoff et al., 2021; Nacson et al., 2023). We add to this literature a two brand-new dimensions: how a concentration coefficient of data distribution affects generalization by EoS and how gradient descent with large step-size adapts to low-dimensional structures in data.
@@
More broaderly, our work is inspired by the seminal work of Zhang et al. (2017) on “rethinking generalization”. Our results provide new theoretical justification that rigorously explains several curious phenomena (such as why real data are harder to overfit than random Gaussian data) reported therein. Compared to other existing work inspired by Zhang et al. (2017), e.g., those that study the implicit bias of gradient descent from various alternative angles (dynamics (Arora et al., 2019; Mei et al., 2019; Jin and Montúfar, 2023), algorithmic stability (Hardt et al., 2016), large-margin (Soudry et al., 2018), benign overfitting (Joshi et al., 2024; Kornowski et al., 2024)), our work has more end-to-end generalization bounds and requires (morally, since the settings are not all compatible) weaker assumptions. On the practical front, we provide new theoretical insight into how “mix-up” data augmentation (Zhang et al., 2018, 2021) and “activation-based pruning” (Hu et al., 2016; Ganguli and Chong, 2024) work. A more detailed discussion of the related work and the implications of our results can be found in Appendix B.
Scope of Analysis. Our theoretical framework is situated in the feature-learning (or “rich”) regime of overparameterized networks. To bypass the direct analysis of complex gradient dynamics, we characterize the properties of the set of dynamically stable solutions reachable by gradient descent. This set is defined by the Below-Edge-of-Stability (BEoS) condition, which holds not only for final solutions but also for network states throughout extended periods of training.
Our bounds therefore quantify a foundational generalization guarantee that is persistently enforced by this stability requirement. While other transient implicit biases might yield even tighter generalization in early training, our work characterizes the fundamental regularization effect stemming from stability. A key advantage of this approach is its width-agnostic nature, allowing our generalization upper bounds to apply to networks of arbitrary finite width. In contrast, our negative results, the lower-bound construction (Theorem 3.6) and the stable interpolators (Theorem 3.7), require a network width $K$ at least on the order of the sample size $n$.
Our theoretical results are derived for two-layer fully-connected ReLU networks, a cornerstone for theoretical analysis. Extending this framework to deeper networks or architectures with specific inductive biases (e.g., CNNs) is a significant undertaking that we leave for future work.
@@ 2 Preliminaries and Notations
This condition applies to any twice-differentiable solution found by GD, even when the optimization process does not converge to a local or global minimum. Moreover, BEoS is empirically verified to hold during both the “progressive sharpening” phase and the subsequent oscillatory phase at the EoS.
@@
Here, $\bm{X}$} is a random vector drawn uniformly at random from the training examples $\{{\bm{x}}_{i}\}_{i=1}^{n}$. Specifically, we may also consider its population level $g_{\mathcal{P}}$ by viewing $\bm{X}$} as a random variable
@@
The proof of this proposition refers to (Liang et al., 2025, Corollary 3.3). The non-parametric characterization of stable minima via bounded weighted variation norm refers to (Liang et al., 2025; Nacson et al., 2023).
@@
3 Main Results
In this section, we present our main theoretical results concerning the properties of stable solutions found by gradient descent. Section 3.1 establishes a generalization bound for data exhibiting intrinsic low-dimensional structure. Section 3.2 then derives a spectrum of generalization bounds for a tunable family of isotropic distributions, which connects the data’s radial mass concentration to generalization performance. Finally, Section 3.3 investigates the behavior of stable solutions in the limiting case where data is supported entirely on the unit sphere. All the detail proofs are deferred to the appendix.
3.1 Provable Adaptation to Intrinsic Low-Dimensionality
We begin with our main positive result, considering the case where data possesses an underlying low-dimensional structure, a common feature of real-world datasets. We show that in this context, the generalization performance of stable networks adapts to this intrinsic dimension rather than the ambient one.
Assumption 3.1 (Mixture of Low-Dimensional Balls).
where $p_{j}>0$ are the mixture probabilities $\mathbb{P}({\bm{x}}\in V_{j})$ satisfying $\sum_{j=1}^{J}p_{j}=1$. Each component distribution $\mathcal{P}_{j}$ is the uniform distribution on the unit disc $\mathbb{B}_{1}^{V_{j}}:=\{{\bm{x}}\in V_{j}:\left\lVert{\bm{x}}\right\rVert_{2}\leq 1\}$. The corresponding labels $y$ are generated from a conditional distribution $\mathcal{P}(y|{\bm{x}})$ and are assumed to be bounded, i.e., $|y|\leq D$ for some constant $D>0$. Similarly, we define $\mathcal{P}_{j}({\bm{x}},y)=\mathcal{P}({\bm{x}},y\mid{\bm{x}}\in V_{j})$.
Under the structural conditions of Assumption 3.1, we establish a generalization bound whose sample complexity depends on the intrinsic dimension $m$.
Theorem 3.2 (Generalization Bound for Mixture Models).
Let the data distribution $\mathcal{P}$ be as defined in Assumption 3.1. Let $\mathcal{D}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$ be a dataset of $n$ i.i.d. samples drawn from $\mathcal{P}$. Then, with probability at least $1-\delta$,
$\sup_{{{\bm{\theta}}}\in{\bm{\Theta}}_{\mathrm{BEoS}}(\eta,\mathcal{D})}\mathrm{Gap}_{\mathcal{P}}(f_{{\bm{\theta}}};\mathcal{D})\;\lessapprox_{d}\;\Big(\frac{1}{\eta}-\frac{1}{2}+4M\Big)^{\frac{m}{m^{2}+4m+3}}\,M^{2}\,J^{\frac{4}{m}}\,n^{-\frac{1}{2m+4}}+M^{2}\,J\,\sqrt{\frac{1}{2n}}.$ (7)
where $M\coloneqq\max\{D,\|f_{\bm{\theta}}|_{\mathbb{B}_{1}^{V}}\|_{L^{\infty}},1\}$ and $\lessapprox_{d}$ hides constants (which could depend on $d$) and logarithmic factors in $J/\delta$ and $n$.
The proof appears in Appendix F. The core strategy is to decompose the problem by analyzing the network’s behavior on each subspace $V_{j}$ individually and then aggregating the results. If we restrict the network to a single $m$-dimensional subspace $V_{j}$, a neuron’s activation is governed not by its full weight vector ${\bm{w}}_{k}$, but solely by $\operatorname{proj}_{V_{j}}{\bm{w}}_{k}$, since the component of ${\bm{w}}_{k}$ that is orthogonal to $V_{j}$ is “invisible to the data on $V_{j}$”. However, a critical question then emerges: we know EoS provides a constraint on the global data-dependent regularity, which involves the full weights $w_{k}$, but how does this global constraint translate to local constraint on each individual subspace? To resolve this question, we prove that the global weight function $g$ dominates the local weight function $g_{j}$ that is only determined by the data points on $V_{j}$ (Lemma F.3). This formally establishes that the implicit regularization is adaptive to the data’s geometry, allowing us to derive a final bound that scales with the intrinsic dimension $m$.

### v2

@@ Abstract
Understanding generalization in overparameterized neural networks hinges on the interplay between the data geometry, neural architecture, and training dynamics. In this paper, we theoretically explore how data geometry controls this implicit bias. This paper presents theoretical results for overparametrized two-layer ReLU networks trained below the edge of stability. First, for data distributions supported on a mixture of low-dimensional balls, we derive generalization bounds that provably adapt to the intrinsic dimension. Second, for a family of isotropic distributions that vary in how strongly probability mass concentrates toward the unit sphere, we derive a spectrum of bounds showing that rates deteriorate as the mass concentrates toward the sphere. These results instantiate a unifying principle: When the data is harder to “shatter” with respect to the activation thresholds of the ReLU neurons, gradient descent tends to learn representations that capture shared patterns and thus finds solutions that generalize well. On the other hand, for data that is easily shattered (e.g., data supported on the sphere) gradient descent favors memorization. Our theoretical results consolidate disparate empirical findings that have appeared in the literature.
@@ 1 Introduction
Contributions. We argue that the effectiveness of this data-dependent regularity is governed by a single geometric quantity, which we call data shatterability: qualitatively, how easily the data distribution can be partitioned into many disjoint small regions by ReLU half-spaces. Informally, the less shatterable the data geometry, the stronger the implicit regularization at the EoS. We make this principle precise and obtain the following results.
@@
  • A Spectrum of Generalization on Isotropic Data. For a one-parameter family of isotropic $\mathrm{Beta}(\alpha)$-radial distributions, we derive generalization upper bounds and matching lower bounds that depend smoothly on the radial concentration parameter $\alpha$ (Theorems 3.4 and 3.5). As $\alpha$ decreases, the probability mass moves towards the boundary and the generalization guarantee degrades. In the limiting case where the support collapses to the unit sphere, we construct perfectly interpolating networks that still satisfy the BEoS stability condition (Theorem 3.6). In particular, the “neural shattering” phenomenon, identified by Liang et al. (2025) for the uniform ball distribution, represents one special point of the broader generalization spectrum we uncover.
  • Provable Adaptation to Low-dimensionality. Under a mixture-of-subspaces assumption where inputs are supported on a union of $m$-dimensional balls in $\mathbb{R}^{d}$ with $m<d$, we prove that all BEoS-stable solutions enjoy a generalization rate $\tilde{O}\big(n^{-1/(2m+4)}\big)$ that depends on the intrinsic dimension $m$ rather than the ambient dimension $d$ (Theorem 3.10). The dependence on the number of mixture components is at most polynomial. Synthetic experiments confirm that gradient descent indeed behaves according to this intrinsic-dimension law.
Taken together, these results identify data shatterability as the key geometric quantity that controls the strength of implicit regularization for gradient descent below the edge of stability.
Technical novelty. Many classical uniform-convergence-based generalization bounds for overparameterized neural networks are distribution-agnostic: one first controls the $L^{\infty}$ metric entropy of a function class, then controls the Rademacher complexity (or derive a uniform concentration inequality via a union bound). This approach is not available in our setting. The data-dependent regularity induced by the EoS condition defines a function class whose $L^{\infty}$ metric entropy is infinite, and our spherical interpolation result shows that nontrivial distribution-agnostic bounds cannot hold. The key observation is that stability-induced regularity is highly inhomogeneous over the input domain: there are regions where the effective regularization is strong (“good” regions) and regions where it is extremely weak (“bad” regions) such that metric entropy explodes.
Our main technical innovation is to bypass global metric entropy control via a half-space-depth quantile partition of the input space. This technique allows us to decouple the analysis based on the strength of regularization: In the “good region”, the implicit regularization is effective, allowing us to enforce strict complexity control, while in the “bad region”, where complexity explodes, we abandon function-space covering arguments. Instead, we control the generalization error by bounding the probability mass of these regions, tying the error directly to the data geometry.
Conceptually, this framework inverts the classical VC dimension perspective. While VC dimension characterizes a model’s active capacity to shatter arbitrary data, our “data shatterability” principle characterizes the feasibility of a specific dataset being shattered by the GD-trained network.
@@
Related work. We build upon a recent line of work (Wu and Su, 2023; Qiao et al., 2024; Liang et al., 2025) that theoretically study the generalization of neural networks in Edge-of-Stability regime (Cohen et al., 2020) from a function space perspective (Mulayoff et al., 2021; Nacson et al., 2023). We add to this literature by abstracting neural-shattering analysis on uniform-ball of Liang et al. (2025) into a depth-based data-shatterability framework and extending it to a broader class of distributions that capture both radial concentration and low-dimensional structure, yielding distribution-dependent upper and lower bounds that make the role of data geometry explicit.
@@
More broadly, our work is inspired by the seminal work of Zhang et al. (2017) on “rethinking generalization”. Our results provide new theoretical justification that rigorously explains several curious phenomena (such as why real data are harder to overfit than random Gaussian data) reported therein. Compared to other existing work inspired by Zhang et al. (2017), e.g., those that study the implicit bias of gradient descent from various alternative angles (dynamics (Arora et al., 2019; Mei et al., 2019; Jin and Montúfar, 2023), algorithmic stability (Hardt et al., 2016), large-margin (Soudry et al., 2018), benign overfitting (Joshi et al., 2024; Kornowski et al., 2024)), our work has more end-to-end generalization bounds and requires (morally, since the settings are not all compatible) weaker assumptions. On the practical front, we provide new theoretical insight into how “mix-up” data augmentation (Zhang et al., 2018; 2021) and “activation-based pruning” (Hu et al., 2016; Ganguli and Chong, 2024) work. A more detailed discussion of the related work and the implications of our results can be found in Appendix B.1.
Scope of analysis. Our theoretical framework is situated in the feature-learning (or “rich”) regime of overparameterized networks. Rather than tracking the full gradient dynamics [^1] Nevertheless, there is informal and heuristic discussion from the perspective of gradient dynamics in Appendix B.2., we focus on the regime in which gradient descent with a large step size operates for long stretches below around the edge-of-stability boundary. We therefore analyze the generalization behavior of all parameter vectors satisfying a Below-Edge-of-Stability (BEoS, see Definition 2.1), without assuming optimality or stationarity. Our bounds hold uniformly over this BEoS region and characterize a baseline form of implicit regularization that is enforced whenever training remains near the edge of stability.
@@ 2 Preliminaries and Notations
This condition applies to any twice-differentiable solution found by GD, even when the optimization process does not converge to a local or global minimum. Moreover, BEoS is empirically verified to hold during both the “progressive sharpening” phase and the subsequent oscillatory phase.
@@
Here, $\bm{X}$} is a random vector drawn uniformly at random from the training examples $\{{\bm{x}}_{i}\}_{i=1}^{n}$. Specifically, we may also consider its population level $g_{\mathcal{P}}$ by viewing $\bm{X}$} as a random variable.
Informally, $g_{\mathcal{D}}({\bm{u}},t)$ measures how hard for GD to place a ReLU ridge with normalized $\bm{u}$} and threshold $t$ under the BEoS constraint. Large $g_{\mathcal{D}}$ corresponds to directions that sense that sense significant data activation, thereby yielding strong gradients and hence strong implicit regularization. A more detailed interpretation from the perspective of gradient dynamics is given in Appendix B.2.
@@
The proof of this proposition refers to (Liang et al., 2025, Corollary 3.3). The non-parametric version of the weighted path norm constrain can be found in (Liang et al., 2025; Nacson et al., 2023).
@@
3 Data Shatterability Principle and Generalization Bounds
This section introduces our notion of data shatterability and explains how it leads to a unified proof strategy for both the generalization upper bounds (Theorem 3.10, Theorem 3.4) and the lower bounds (Theorem 3.5, Theorem 3.6).
We start with the classical notion of half-space depth (also known as Tukey depth).
Definition 3.1.
Given a distribution $\mathcal{P}_{X}$ on $\mathbb{R}^{d}$. For any ${\bm{x}}\in\mathbb{R}^{d}$, its population half-space depth $\mathrm{depth}({\bm{x}},\mathcal{P}_{X})=\inf_{{\bm{u}}\in\mathbb{S}^{d-1}}\mathbb{P}_{{\bm{X}}\sim\mathcal{P}_{X}}\big({\bm{u}}^{\mathsf{T}}\left({\bm{X}}-{\bm{x}}\right)\geq 0)$. On the other hand, given a point cloud $\mathcal{X}=\{{\bm{x}}_{1},\cdots,{\bm{x}}_{n}\}\subset\mathbb{R}^{d}$ that is also represented by the empirical distribution $\mathcal{P}(\mathcal{X})$ that assigns mass $1/n$ to each point, the empirical half-space depth $\mathrm{depth}({\bm{x}},\mathcal{X})=\inf_{{\bm{u}}\in\mathbb{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\left\{{\bm{u}}^{\mathsf{T}}({\bm{x}}_{i}-{\bm{x}})\geq 0\right\}$.
Half-space depth measures the centrality of a point $\bm{x}$} by finding the minimum data mass (either population or empirical) on one side of any hyperplane passing through it (Tukey, 1975). For any $T\in[0,\frac{1}{2}]$, we define $T$-deep region $\Omega_{T}(\mathcal{P}_{X})=\{{\bm{x}}\in\mathbb{R}^{d}\mid\mathrm{depth}({\bm{x}},\mathcal{P}_{X})\geq T\}$ as an upper level set of the half-space depth function (we define $\Omega_{T}(\mathcal{X})$ similarly).
Now let $\mathcal{X}=\{{\bm{x}}_{1},\ldots,{\bm{x}}_{n}\}$ be the point cloud of inputs in the dataset $\mathcal{D}$. The key geometric observation is that, for a ReLU neuron to create nonlinearity inside a $T$-deep region $\Omega_{T}(\mathcal{X})$, its activation boundary must intersect that region. Formally, let $f_{{\bm{\theta}}}$ be a network of the form (1) and define $N_{T}:=\Big\{k\,\Big|\,\{{\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}-b_{k}=0\}\cap\Omega_{T}(\mathcal{X})\neq\emptyset\Big\}.$ The neurons with $k\notin N_{T}$ are either always active or always inactive on $\Omega_{T}(\mathcal{X})$, hence contribute only an affine function on that region. Therefore there exists an affine function ${\bm{x}}\mapsto{\bm{c}}^{\mathsf{T}}{\bm{x}}+b$ (which absorbs all neurons with $k\notin N_{T}$) such that $f_{{\bm{\theta}}}({\bm{x}})=\sum_{k\in N_{T}}v_{k}\phi({\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}-b_{k})+{\bm{c}}^{\mathsf{T}}{\bm{x}}+b$, $\forall\,{\bm{x}}\in\Omega_{T}(\mathcal{X})$.
By the definition of half-space depth, every hyperplane passing through a point in $\Omega_{T}(\mathcal{X})$ leaves at least a $T$-fraction of the data on each side. In particular, for any neuron whose boundary intersects $\Omega_{T}(\mathcal{X})$, its activation event has probability at least $T$ (population or empirical, depending on the context). This gives a positive lower bound on the data-dependent weight function $g$. We denote
$g_{\mathcal{D}}^{\min}(T):=\inf_{\{{\bm{u}}^{\mathsf{T}}{\bm{x}}-t=0\}\cap\Omega_{T}(\mathcal{X})\neq\emptyset}g_{\mathcal{D}}({\bm{u}},t)>0,\qquad g_{\mathcal{P}}^{\min}(T):=\inf_{\{{\bm{u}}^{\mathsf{T}}{\bm{x}}-t=0\}\cap\Omega_{T}(\mathcal{P}_{X})\neq\emptyset}g_{\mathcal{P}}({\bm{u}},t).$
Thus, the neurons that generate nonlinearity on $\Omega_{T}(\mathcal{X})$ are effectively regularized by the BEoS condition, with strength controlled from below by $g_{\mathcal{D}}^{\min}(T)$ (and $g_{\mathcal{P}}^{\min}(T)$ at the population level).
On the $T$-deep region, only neurons in $N_{T}$ contribute nonlinearity, and for these we have $g_{\mathcal{D}}({\bm{u}}_{k},t_{k})\geq g_{\mathcal{D}}^{\min}(T)$. Hence their unweighted path norm satisfies $\sum_{k\in N_{T}}|v_{k}|\,\|{\bm{w}}_{k}\|_{2}\leq(g_{\mathcal{D}}^{\min}(T))^{-1}|\|f_{{\bm{\theta}}}\|_{\mathrm{path},g_{\mathcal{D}}}$, so the restriction of $f_{{\bm{\theta}}}$ to $\Omega_{T}(\mathcal{X})$ lies in a standard (unweighted) path-norm ball with radius proportional to $(g_{\mathcal{D}}^{\min}(T))^{-1}$. Generalization bounds for such unweighted path-norm classes on bounded inputs are known (Parhi and Nowak, 2023; Neyshabur et al., 2015) and yield the $T$-deep term in (6). Outside the $T$-deep region, BEoS provides only weak control, so we bound the contribution by the worst-case $L^{\infty}$ amplitude times the probability mass of the shallow region. Finally, since $g_{\mathcal{D}}^{\min}(T)$ and $\Omega_{T}(\mathcal{X})$ are random, we replace them by their population counterparts $g_{\mathcal{P}}^{\min}(T)$ and $\Omega_{T}(\mathcal{P}_{X})$ and control the discrepancy using empirical process tools (Appendix E.2). This yields the following generic decomposition, holding with high probability
$\sup_{{\bm{\theta}}\in{{\bm{\Theta}}_{\mathrm{BEoS}}}(\eta,\mathcal{D})}\mathrm{Gap}(f_{{\bm{\theta}}},\mathcal{D})\;\leq\;\underbrace{\tilde{O}\Big(\mathbb{P}_{{\bm{X}}}({\bm{X}}\notin\Omega_{T}(\mathcal{P}_{X}))\Big)}_{\text{shallow region}}\;+\;\underbrace{\tilde{O}\Big(g_{\mathcal{P}}^{\min}(T)^{-\frac{d}{2d+3}}\,n^{-\frac{d+3}{4d+6}}\Big)}_{\text{$T$-deep region}},$ (6)
where $\tilde{O}$ hides logarithmic factors and universal constants. This inequality is the main technical bridge between data geometry and the behavior of BEoS solutions. Now we instantiate it for isotropic and anisotropic distributions to establish the corresponding generalization bounds.